# NuSegDG: Integration of Heterogeneous Space and Gaussian Kernel for Domain-Generalized Nuclei Segmentation

### v2

@@ 3.4 Two-Stage Mask Decoder
In the last decade, U-shape hierarchical decoders [7, 34, 9] have been widely used for the prediction of nuclei semantic and instance segmentation masks. For the latter, previous methods usually utilized morphological post-processing methods to detect each cell based on the generated nuclear proxy maps. However, such operations require laboriously manual parameter adjustment when facing different nuclei domains, degrading the generalization capabilities of models. On the other hand, current medical SAMs [15, 16, 13] adopted a sequential inference algorithm to recognize each target object in images, so they are time-consuming for nuclei instance segmentation tasks involving a large number of cells. To address this issue, we propose the TSM-Decoder that improves the efficiency of producing instance maps by focusing on the prediction of precise semantic segmentation masks. Specifically, we first create trainable query embeddings $q\in\mathbb{R}^{C\times 256}$ to save the decoding information. Different from SAM [10], $C$ represents the number of prediction categories instead of multi-layer masks as histopathology images may include different types of nuclei. Then, we concatenate position prompts $\mathcal{P}_{pos}$ with $q$ and perform a self-attention operation as:
@@
where $\Psi$ is positional encodings. Similar to SAM [10], we iterate this operation twice for sufficient updation. Finally, we predict the semantic segmentation mask $\hat{\mathcal{Y}}_{k}\in\mathbb{R}^{H\times W}$ by:
@@
Table 1: Domain-generalized comparison with state-of-the-arts on nuclei semantic segmentation.
Datasets $\mathcal{T}=\mathcal{S}_{1}$ $\mathcal{T}=\mathcal{S}_{2}$ $\mathcal{T}=\mathcal{S}_{3}$ $\mathcal{T}=\mathcal{S}_{4}$
Methods Dice mIoU F1 HD Dice mIoU F1 HD Dice mIoU F1 HD Dice mIoU F1 HD
U-Net [20] 21.88 15.31 26.52 355.19 56.01 41.66 65.81 77.08 25.73 16.87 30.73 399.99 65.49 49.89 68.52 99.06
U-Net++ [21] 25.99 17.53 31.55 372.39 59.64 46.03 67.26 86.35 20.14 12.50 22.50 418.79 66.54 51.04 68.80 99.63
AttUNet [22] 26.66 18.30 32.42 372.44 58.79 45.24 66.79 92.43 27.70 17.98 31.74 375.44 66.13 50.55 68.46 101.81
DCSAU-Net [23] 43.42 30.73 47.49 246.01 67.17 52.72 71.83 78.62 33.66 22.34 37.27 307.45 71.58 56.47 73.25 94.36
TransUNet [28] 64.41 52.85 70.93 151.67 77.51 63.54 78.11 64.64 73.41 60.26 76.53 229.66 74.17 59.43 75.27 95.42
ACC-UNet [50] 29.11 20.27 35.12 381.36 65.75 51.39 70.54 77.79 32.51 21.29 35.35 312.20 69.63 54.24 70.94 98.92
nnU-Net [51] 23.71 16.34 28.48 360.33 62.43 48.37 69.79 81.52 31.45 20.75 34.70 342.90 68.88 53.50 70.67 99.56
U-mamba [29] 12.01 9.05 14.89 392.97 51.68 39.28 62.53 101.20 19.39 11.89 21.50 408.82 58.79 45.24 66.79 92.43
FedDG [38] 62.79 49.43 69.99 266.36 77.03 62.90 78.01 70.12 71.19 57.65 74.47 242.18 71.85 56.64 72.83 99.19
DCAC [37] 50.29 38.01 55.50 162.13 68.50 53.71 72.48 73.95 29.96 19.48 33.36 322.49 69.71 54.36 71.15 100.21
SAM [10] 66.59 55.29 72.37 252.80 76.84 62.63 77.38 66.82 76.48 63.16 78.34 144.39 75.56 61.25 76.14 75.61
Med-SA [43] 76.94 66.21 80.73 125.24 79.55 66.25 80.00 65.19 78.43 66.26 81.02 127.02 80.54 67.66 80.75 80.75
SAMed [47] 76.79 65.70 80.27 137.15 78.86 65.35 79.51 66.20 78.48 65.68 80.06 137.82 78.77 65.29 79.08 69.09
SAMUS [44] 78.51 68.09 82.14 107.32 80.25 67.16 80.68 63.27 80.91 68.69 82.40 82.33 80.72 67.98 80.97 91.39
SAM-CL [41] 73.48 63.05 78.65 157.08 77.51 63.54 78.11 64.64 76.99 63.94 79.24 119.73 78.65 65.12 78.95 74.44
SAC [52] 69.11 58.14 74.60 245.04 77.80 63.89 78.52 66.76 76.09 63.17 78.49 212.19 75.32 60.86 75.74 90.28
LeSAM [53] 71.80 61.45 77.55 207.65 77.83 63.95 78.58 67.65 76.48 63.78 78.80 205.46 74.72 60.02 75.06 91.03
H-SAM [54] 76.79 65.70 80.73 165.25 78.00 64.17 78.65 66.07 76.33 64.03 78.89 215.47 77.04 63.00 77.40 89.83
@@
Table 2: Adaptability Comparison with state-of-the-arts on nuclei semantic segmentation.
Datasets $\mathcal{S}_{1}$ $\mathcal{S}_{2}$ $\mathcal{S}_{3}$ $\mathcal{S}_{4}$
Methods Dice mIoU F1 HD Dice mIoU F1 HD Dice mIoU F1 HD Dice mIoU F1 HD
U-Net [20] 90.42 83.22 91.07 137.11 71.41 56.17 72.06 78.13 72.27 57.38 72.97 123.13 81.02 68.52 81.19 90.01
U-Net++ [21] 90.85 83.83 91.33 117.58 75.72 61.05 76.15 73.38 64.66 49.61 67.56 148.37 81.80 69.62 91.97 80.32
AttUNet [22] 91.01 84.13 91.41 112.98 75.81 61.23 76.13 74.32 75.98 61.81 77.28 131.51 81.34 68.98 81.69 80.75
DCSAU-Net [23] 91.74 85.15 92.04 127.16 75.19 60.38 75.58 77.22 78.33 64.61 78.99 103.61 80.90 68.42 81.24 83.89
TransUNet [28] 91.41 84.65 91.75 132.94 76.30 61.83 76.70 77.33 76.59 62.38 77.51 109.36 82.49 70.47 82.60 78.14
ACC-UNet [50] 90.95 83.96 91.38 119.39 77.90 63.93 78.32 74.79 66.84 52.32 70.29 147.97 82.13 69.91 82.23 77.89
nnU-Net [51] 90.11 82.45 90.57 138.83 80.84 67.91 81.03 73.04 84.32 72.91 84.43 117.51 81.37 68.73 81.63 83.01
U-mamba [29] 89.96 82.58 90.66 127.03 77.38 63.23 77.74 73.10 63.62 47.81 65.44 167.67 82.43 70.37 82.57 82.42
FedDG [38] 90.48 83.25 91.11 134.97 74.57 59.77 75.16 76.74 74.88 60.45 76.57 110.62 81.51 69.17 81.93 86.56
DCAC [37] 91.22 84.35 91.60 112.82 74.89 60.11 75.32 74.31 63.79 49.75 68.14 152.78 81.53 69.19 81.65 71.69
SAM [10] 89.78 82.14 90.33 138.58 76.87 62.63 77.23 78.83 79.28 65.91 79.90 153.47 77.75 64.43 78.39 101.65
Med-SA [43] 91.72 85.12 92.08 69.42 81.32 68.61 81.57 74.59 83.47 71.89 83.63 86.62 83.32 71.59 83.37 77.74
SAMed [47] 91.32 84.47 91.63 71.67 80.06 66.97 80.46 77.14 81.75 69.20 81.78 101.22 82.23 70.05 82.32 76.68
SAMUS [44] 92.07 85.67 92.35 115.18 83.26 71.39 83.33 71.14 84.67 73.45 84.72 114.98 83.40 71.75 83.70 79.15
SAM-CL [41] 91.81 85.29 92.12 89.48 81.17 68.41 81.34 73.57 82.95 70.90 83.02 135.31 82.51 70.46 82.86 90.85
SAC [52] 91.79 85.16 92.10 123.01 81.06 68.25 81.21 72.68 83.51 71.74 83.72 119.45 81.80 69.36 81.96 80.83
LeSAM [53] 91.71 85.08 92.00 128.00 79.67 66.34 79.85 74.36 82.43 70.15 82.57 137.49 80.38 67.56 80.90 90.91
H-SAM [54] 92.01 85.53 92.27 129.34 81.45 68.81 81.55 71.99 83.40 71.60 83.51 118.60 81.65 69.20 82.09 82.24
@@ 4.1.1 Datasets
To validate the effectiveness of the proposed NuSegDG, we collect DSB-2018 [3], MoNuSeg-2018 [4], TNBC [5] and CryoNuSeg [6] datasets to perform comprehensive comparisons for domain generalization. We denote these four nuclei datasets with source domains $\mathcal{S}_{1}$, $\mathcal{S}_{2}$, $\mathcal{S}_{3}$ and $\mathcal{S}_{4}$, respectively. The details are as follows.
@@
DSB-2018 [3] dataset includes 670 nuclei images captured using fluorescence microscopy, offering a range of staining methods including DAPI, Hoechst, Hematoxylin and Eosin (H&E). These images are annotated with nuclear masks to facilitate segmentation tasks and vary in size.
@@
MoNuSeg-2018 [4] dataset consists of 51 H&E stained histopathology images from various organs, including breast, liver, kidney, prostate, bladder, colon, and stomach. Each image measures $1000\times 1000$ pixels, captured at 40×40$\times$40 × magnification.
@@
TNBC [5] dataset comprises nuclei images stained with Hematoxylin and Eosin (H&E), sourced from breast cancer patients. This dataset includes 50 images with a resolution of $512\times 512$ pixels, captured at 40×40$\times$40 × magnification.
@@
CryoNuSeg [6] dataset contains H&E stained tissues from 10 different organs, providing 30 images of $512\times 512$ pixels, captured at 40×40$\times$40 × magnification. The diversity of tissue types offers a comprehensive resource for evaluating the robustness of segmentation methods.

### v3

@@ 3.4 Two-Stage Mask Decoder
In the last decade, U-shape hierarchical decoders [graham2019hover, ilyas2022tsfd, horst2024cellvit] have been widely used for the prediction of nuclei semantic and instance segmentation masks. For the latter, previous methods usually utilized morphological post-processing methods to detect each cell based on the generated nuclear proxy maps. However, such operations require laboriously manual parameter adjustment when facing different nuclei domains, degrading the generalization capabilities of models. On the other hand, current medical SAMs [ma2024segment, huang2024segment, shui2023unleashing] adopted a sequential inference algorithm to recognize each target object in images, so they are time-consuming for nuclei instance segmentation tasks involving a large number of cells. To address this issue, we propose the TSM-Decoder that improves the efficiency of producing instance maps by focusing on the prediction of precise semantic segmentation masks. Specifically, we first create trainable query embeddings $q\in\mathbb{R}^{C\times 256}$ to save the decoding information. Different from SAM [kirillov2023segment], $C$ represents the number of prediction categories instead of multi-layer masks as histopathology images may include different types of nuclei. Then, we concatenate position prompts $\mathcal{P}_{pos}$ with $q$ and perform a self-attention operation as:
@@
where $\Psi$ is positional encodings. Similar to SAM [kirillov2023segment], we iterate this operation twice for sufficient updation. Finally, we predict the semantic segmentation mask $\hat{\mathcal{Y}}_{k}\in\mathbb{R}^{H\times W}$ by:
@@
Table 1: Comparison with state-of-the-arts on nuclei semantic segmentation (Source Domain Generalization).
Datasets Manual $\mathcal{S}_{1}$ $\mathcal{S}_{2}$ $\mathcal{S}_{3}$ $\mathcal{S}_{4}$
Methods Prompt Dice mIoU F1 HD Dice mIoU F1 HD Dice mIoU F1 HD Dice mIoU F1 HD
U-Net [ronneberger2015u] 90.42 83.22 91.07 137.11 71.41 56.17 72.06 78.13 72.27 57.38 72.97 123.13 81.02 68.52 81.19 90.01
U-Net++ [zhou2019unet++] 90.85 83.83 91.33 117.58 75.72 61.05 76.15 73.38 64.66 49.61 67.56 148.37 81.80 69.62 91.97 80.32
AttUNet [schlemper2019attention] 91.01 84.13 91.41 112.98 75.81 61.23 76.13 74.32 75.98 61.81 77.28 131.51 81.34 68.98 81.69 80.75
DCSAU-Net [xu2023dcsau] 91.74 85.15 92.04 127.16 75.19 60.38 75.58 77.22 78.33 64.61 78.99 103.61 80.90 68.42 81.24 83.89
TransUNet [chen2024transunet] 91.41 84.65 91.75 132.94 76.30 61.83 76.70 77.33 76.59 62.38 77.51 109.36 82.49 70.47 82.60 78.14
ACC-UNet [ibtehaz2023acc] 90.95 83.96 91.38 119.39 77.90 63.93 78.32 74.79 66.84 52.32 70.29 147.97 82.13 69.91 82.23 77.89
nnU-Net [isensee2021nnu] 90.11 82.45 90.57 138.83 80.84 67.91 81.03 73.04 84.32 72.91 84.43 117.51 81.37 68.73 81.63 83.01
U-mamba [ma2024u] 89.96 82.58 90.66 127.03 77.38 63.23 77.74 73.10 63.62 47.81 65.44 167.67 82.43 70.37 82.57 82.42
FedDG [liu2021feddg] 90.48 83.25 91.11 134.97 74.57 59.77 75.16 76.74 74.88 60.45 76.57 110.62 81.51 69.17 81.93 86.56
DCAC [hu2022domain] 91.22 84.35 91.60 112.82 74.89 60.11 75.32 74.31 63.79 49.75 68.14 152.78 81.53 69.19 81.65 71.69
SAC [na2024segment] 91.79 85.16 92.10 123.01 81.06 68.25 81.21 72.68 83.51 71.74 83.72 119.45 81.80 69.36 81.96 80.83
H-SAM [cheng2024unleashing] 92.01 85.53 92.27 129.34 81.45 68.81 81.55 71.99 83.40 71.60 83.51 118.60 81.65 69.20 82.09 82.24
SAM [kirillov2023segment] 89.78 82.14 90.33 138.58 76.87 62.63 77.23 78.83 79.28 65.91 79.90 153.47 77.75 64.43 78.39 101.65
Med-SA [wu2023medical] 91.72 85.12 92.08 69.42 81.32 68.61 81.57 74.59 83.47 71.89 83.63 86.62 83.32 71.59 83.37 77.74
SAMed [zhang2023customized] 91.32 84.47 91.63 71.67 80.06 66.97 80.46 77.14 81.75 69.20 81.78 101.22 82.23 70.05 82.32 76.68
LeSAM [gu2024lesam] 91.71 85.08 92.00 128.00 79.67 66.34 79.85 74.36 82.43 70.15 82.57 137.49 80.38 67.56 80.90 90.91
SAMUS [lin2023samus] 92.07 85.67 92.35 115.18 83.26 71.39 83.33 71.14 84.67 73.45 84.72 114.98 83.40 71.75 83.70 79.15
SAM-CL [zhong2024convolution] 91.81 85.29 92.12 89.48 81.17 68.41 81.34 73.57 82.95 70.90 83.02 135.31 82.51 70.46 82.86 90.85
Table 2: Comparison with state-of-the-arts on nuclei semantic segmentation (Target Domain Generalization).
Datasets Manual $\mathcal{T}=\mathcal{S}_{1}$ $\mathcal{T}=\mathcal{S}_{2}$ $\mathcal{T}=\mathcal{S}_{3}$ $\mathcal{T}=\mathcal{S}_{4}$
Methods Prompt Dice mIoU F1 HD Dice mIoU F1 HD Dice mIoU F1 HD Dice mIoU F1 HD
U-Net [ronneberger2015u] 21.88 15.31 26.52 355.19 56.01 41.66 65.81 77.08 25.73 16.87 30.73 399.99 65.49 49.89 68.52 99.06
U-Net++ [zhou2019unet++] 25.99 17.53 31.55 372.39 59.64 46.03 67.26 86.35 20.14 12.50 22.50 418.79 66.54 51.04 68.80 99.63
AttUNet [schlemper2019attention] 26.66 18.30 32.42 372.44 58.79 45.24 66.79 92.43 27.70 17.98 31.74 375.44 66.13 50.55 68.46 101.81
DCSAU-Net [xu2023dcsau] 43.42 30.73 47.49 246.01 67.17 52.72 71.83 78.62 33.66 22.34 37.27 307.45 71.58 56.47 73.25 94.36
TransUNet [chen2024transunet] 64.41 52.85 70.93 151.67 77.51 63.54 78.11 64.64 73.41 60.26 76.53 229.66 74.17 59.43 75.27 95.42
ACC-UNet [ibtehaz2023acc] 29.11 20.27 35.12 381.36 65.75 51.39 70.54 77.79 32.51 21.29 35.35 312.20 69.63 54.24 70.94 98.92
nnU-Net [isensee2021nnu] 23.71 16.34 28.48 360.33 62.43 48.37 69.79 81.52 31.45 20.75 34.70 342.90 68.88 53.50 70.67 99.56
U-mamba [ma2024u] 12.01 9.05 14.89 392.97 51.68 39.28 62.53 101.20 19.39 11.89 21.50 408.82 58.79 45.24 66.79 92.43
FedDG [liu2021feddg] 62.79 49.43 69.99 266.36 77.03 62.90 78.01 70.12 71.19 57.65 74.47 242.18 71.85 56.64 72.83 99.19
DCAC [hu2022domain] 50.29 38.01 55.50 162.13 68.50 53.71 72.48 73.95 29.96 19.48 33.36 322.49 69.71 54.36 71.15 100.21
SAC [na2024segment] 69.11 58.14 74.60 245.04 77.80 63.89 78.52 66.76 76.09 63.17 78.49 212.19 75.32 60.86 75.74 90.28
H-SAM [cheng2024unleashing] 76.92 66.15 80.67 165.25 78.00 64.17 78.65 66.07 76.33 64.03 78.89 215.47 77.04 63.00 77.40 89.83
SAM [kirillov2023segment] 66.59 55.29 72.37 252.80 76.84 62.63 77.38 66.82 76.48 63.16 78.34 144.39 75.56 61.25 76.14 75.61
Med-SA [wu2023medical] 76.94 66.21 80.73 125.24 79.55 66.25 80.00 65.19 78.43 66.26 81.02 127.02 80.54 67.66 80.75 80.75
SAMed [zhang2023customized] 76.79 65.70 80.27 137.15 78.86 65.35 79.51 66.20 78.48 65.68 80.06 137.82 78.77 65.29 79.08 69.09
LeSAM [gu2024lesam] 71.80 61.45 77.55 207.65 77.83 63.95 78.58 67.65 76.48 63.78 78.80 205.46 74.72 60.02 75.06 91.03
SAMUS [lin2023samus] 78.51 68.09 82.14 107.32 80.25 67.16 80.68 63.27 80.91 68.69 82.40 82.33 80.72 67.98 80.97 91.39
SAM-CL [zhong2024convolution] 73.48 63.05 78.65 157.08 77.51 63.54 78.11 64.64 76.99 63.94 79.24 119.73 78.65 65.12 78.95 74.44
@@ 4.1.1 Datasets
To validate the effectiveness of the proposed NuSegDG, we collect DSB-2018 [caicedo2019nucleus], MoNuSeg-2018 [kumar2017dataset], TNBC [naylor2018segmentation] and CryoNuSeg [mahbod2021cryonuseg] datasets to perform comprehensive comparisons for domain generalization. We denote these four nuclei datasets with source domains $\mathcal{S}_{1}$, $\mathcal{S}_{2}$, $\mathcal{S}_{3}$ and $\mathcal{S}_{4}$, respectively. The details are as follows.
@@
DSB-2018 [caicedo2019nucleus] dataset includes 670 nuclei images captured using fluorescence microscopy, offering a range of staining methods including DAPI, Hoechst, hematoxylin and eosin. These images are annotated with nuclear masks to facilitate segmentation tasks and vary in size.
@@
MoNuSeg-2018 [kumar2017dataset] dataset consists of 51 stained histopathology images from various organs, including breast, liver, kidney, prostate, bladder, colon, and stomach. Each image measures $1000\times 1000$ pixels, captured at 40×40$\times$40 × magnification.
@@
TNBC [naylor2018segmentation] dataset comprises nuclei images stained with hematoxylin and eosin, sourced from breast cancer patients. This dataset includes 50 images with a resolution of $512\times 512$ pixels, captured at 40×40$\times$40 × magnification.
@@
CryoNuSeg [mahbod2021cryonuseg] dataset contains stained tissues from 10 different organs, providing 30 images of $512\times 512$ pixels, captured at 40×40$\times$40 × magnification. The diversity of tissue types offers a comprehensive resource for evaluating the robustness of segmentation methods.
Table 3: Comparison with state-of-the-arts on nuclei instance segmentation (Source Domain Generalization).
Datasets Manual $\mathcal{S}_{1}$ $\mathcal{S}_{2}$ $\mathcal{S}_{3}$ $\mathcal{S}_{4}$
Methods Prompt AJI DQ SQ PQ AJI DQ SQ PQ AJI DQ SQ PQ AJI DQ SQ PQ
U-Net [ronneberger2015u] 63.49 74.81 81.72 61.45 50.27 61.75 74.74 46.23 51.34 60.73 75.32 45.88 46.43 55.90 75.80 42.38
Mask-RCNN [he2017mask] 63.32 75.03 81.05 61.44 45.32 55.98 74.16 41.61 43.84 54.79 74.71 41.35 46.72 56.16 76.25 42.87
StarDIST [caicedo2019nucleus] 63.38 74.78 80.51 60.90 54.95 68.24 74.36 50.87 45.26 56.15 76.11 42.76 46.57 53.82 74.84 40.29
Hover-Net [graham2019hover] 61.59 61.04 79.48 50.03 54.97 71.29 75.68 54.02 25.10 25.14 69.16 17.75 36.35 34.98 71.59 25.21
TSFD-Net [ilyas2022tsfd] 62.25 72.67 80.14 59.06 54.16 67.98 74.55 50.78 42.93 53.77 75.97 40.85 47.64 58.15 75.20 43.79
CellPose [stringer2021cellpose] 66.77 80.21 82.54 66.93 20.62 27.82 74.08 20.72 45.21 62.82 76.88 48.40 36.46 45.90 74.92 34.61
CPP-Net [chen2023cpp] 63.60 74.72 81.51 61.48 52.29 66.21 73.94 49.03 56.00 68.20 77.75 53.05 47.36 56.03 76.00 43.21
CellViT [horst2024cellvit] 60.51 73.92 84.10 63.25 57.91 77.35 77.19 60.54 53.16 68.72 77.65 54.83 40.87 58.13 76.16 44.38
PromptNucSeg [shui2023unleashing] 74.26 85.13 82.25 70.31 64.53 79.03 76.32 60.41 61.94 72.44 77.46 56.98 52.31 62.11 73.23 45.71
SAM [kirillov2023segment] 73.32 80.21 83.08 67.16 55.51 69.70 75.24 52.51 55.16 54.44 77.74 42.45 41.66 50.71 75.71 38.77
Med-SA [wu2023medical] 74.17 84.89 81.95 70.03 64.53 78.64 75.45 59.42 64.59 67.92 77.11 52.89 50.35 59.52 73.80 44.26
SAMed [zhang2023customized] 71.95 82.30 80.04 66.42 62.62 75.52 75.70 57.29 63.30 65.15 75.09 49.09 49.14 56.94 73.94 42.42
SAMUS [lin2023samus] 73.70 86.97 80.88 70.72 67.98 83.09 76.87 63.93 63.34 76.56 77.23 60.84 51.25 63.53 75.79 48.35
SAM-CL [zhong2024convolution] 73.20 86.40 80.89 70.37 63.60 78.49 74.84 58.84 60.11 67.01 77.19 51.90 50.42 61.28 74.58 45.83
Table 4: Comparison with state-of-the-arts on nuclei instance segmentation (Target Domain Generalization).
Datasets Manual $\mathcal{T}=\mathcal{S}_{1}$ $\mathcal{T}=\mathcal{S}_{2}$ $\mathcal{T}=\mathcal{S}_{3}$ $\mathcal{T}=\mathcal{S}_{4}$
Methods Prompt AJI DQ SQ PQ AJI DQ SQ PQ AJI DQ SQ PQ AJI DQ SQ PQ
U-Net [ronneberger2015u] 6.35 4.94 10.74 3.41 31.13 28.70 66.70 19.86 14.24 16.19 48.22 10.88 42.35 47.64 72.10 34.51
Mask-RCNN [he2017mask] 7.73 7.26 12.82 5.08 30.06 30.71 66.49 21.74 11.30 14.05 48.53 9.37 43.57 48.98 71.73 35.30
StarDIST [schmidt2018cell] 8.17 6.55 11.90 4.52 35.58 39.40 70.08 28.49 18.66 24.19 56.50 16.53 43.81 49.61 71.79 35.71
Hover-Net [graham2019hover] 5.42 2.52 16.37 1.75 42.07 53.42 72.96 39.43 23.11 25.50 65.28 17.73 41.10 40.63 64.17 30.48
TSFD-Net [ilyas2022tsfd] 6.03 5.23 11.53 3.68 41.37 49.92 71.41 35.87 12.48 17.02 51.67 11.36 37.95 45.16 71.56 32.50
CellPose [stringer2021cellpose] 15.75 19.00 36.72 13.98 20.56 27.47 73.04 20.18 38.70 54.55 71.16 42.53 41.84 54.39 72.58 41.27
CPP-Net [chen2023cpp] 8.63 6.58 12.04 4.53 41.07 47.81 72.39 34.93 17.42 21.94 60.86 14.66 41.09 47.78 72.33 34.72
CellViT [horst2024cellvit] 4.39 4.95 12.00 3.66 48.44 63.31 73.85 49.17 47.30 64.00 64.76 45.53 43.94 57.23 67.57 42.36
PromptNucSeg [shui2023unleashing] 50.89 48.66 70.44 38.87 48.82 59.06 72.24 43.03 54.14 57.31 72.79 43.78 45.49 44.44 71.73 32.10
SAM [kirillov2023segment] 41.95 39.24 63.75 30.00 47.42 55.73 71.85 40.41 53.92 57.56 72.24 43.79 41.00 48.43 73.03 35.62
Med-SA [wu2023medical] 57.64 55.06 74.90 44.42 50.59 62.06 74.10 46.23 56.60 63.26 72.98 47.94 51.40 61.72 73.50 45.11
SAMed [zhang2023customized] 51.15 46.98 69.71 36.92 49.47 61.00 73.58 45.10 54.59 59.69 73.20 44.14 49.34 59.31 73.33 43.65
SAMUS [lin2023samus] 60.98 61.41 77.80 50.17 50.62 62.17 74.11 46.32 57.27 67.65 72.60 50.32 51.34 62.07 73.40 45.68
SAM-CL [zhong2024convolution] 52.76 51.63 71.08 41.17 47.84 57.26 72.99 42.00 55.29 60.79 73.49 46.30 49.36 59.26 73.59 43.78